Transactions for the month of June were:

                                   Purchases                                                       Sales                

  Quantity Price Q. Price

June 1       (balance) 2,200 @ $3.00                   June 2         900 @ $5.50

                                  3                        3,300 @   3.20                            6      2,400 @   5.50

                                  7                        1,800 @   3.30                            9      1,500 @   5.50

                                15                        2,700 @   3.50                          10         600 @   6.00

                                22                           750 @   3.80                          18      2,100 @   6.00

                                                                                                              25         300 @   6.00

  1. Assuming that periodic inventory, the ending inventory and COGS on a LIFO basis would be (in $)

  1. Assuming that perpetual inventory records are kept, the ending inventory and COGS on a LIFO basis would be (in $)
Answer #1

>> Cost of Goods available for sale =(2200 *3 ) + ( 3300 * 3.2 ) + ( 1800 * 3.3 ) + ( 2700 * 3.5 ) + ( 750 * 3.8 ) = $ 35400.

>> Closing inventory = 10750 - 7800 = 2950 Units.

LIFO Method Periodic Method.

>> Closing Inventory = ( 2200 * 3 ) + ( 750 * 3.2 ) = $ 9000.

>> Cost of Goods sold = $ 35400 - $ 9000 = $ 26400.

LIFO Method Perpetual Method.

>> Closing Inventory = ( 450 * 3.8 ) + ( 600 * 3.5 ) + ( 600 * 3.2 ) + ( 1300 * 3 ) = $ 9630.

>> Cost of Goods sold = $ 35400 - $ 9630 = $ 25770.